The position vector r of a moving particle at time t after the start of the motion is given by

r =5(1+4t)i+5(19+2t−t2)j.

a) Find the initital velocity of the particle.

b) At time t = T, the particle is moving at right angles to its initial direction of motion. Find the value of T and the distance of the particle from its initial position at this time.


1
Expert's answer
2022-01-10T14:05:55-0500

The position vector "r" of a moving particle at time t after the start of the motion is given by

"r =5(1+4t)i+5(19+2t\u2212t^2)j."


a) It follows that the velocity of the particle is given by "v=r'_t =20i+5(2\u22122t)j." Therefore, the initital velocity of the particle is "v_0=20i+10j."


b) Since at time "t = T," the particle is moving at right angles to its initial direction of motion, it folows that the dot product "r(0)\cdot r(T)=0," and hence "5\cdot5(1+4T)+95\cdot 5(19+2T−T^2)=0." Therefore, "1+4T+19(19+2T\u2212T^2)=0," and thus "1+4T+361+38T\u221219T^2=0." It follows that "19T^2-42T-362=0," and "T>0" implies "T=\frac{42+\sqrt{42^2+4\cdot 19\cdot 362}}{38}=\frac{42+\sqrt{29,276}}{38}\approx 5.6."


The distance of the particle from its initial position to the position at time "T" is equal

"5\sqrt{(4T)^2+(2T-T^2)^2}
=5T\sqrt{16+(2-T)^2}
=5T\sqrt{20-4T+T^2}
\approx 150.7."